1 The first section

hello a weird reference a second reference tug to the TEX website which when applied as follows:

2 Random Stuff

This is very dependent on ‘??’ (see p.??).

2.1 url

We can use the \url command to do this as well like so: http://www.tug.org. The \nolinkurl{} is similar but non-functional: http://www.tug.org.

   Now this link http://latex2rtf.sourceforge.net/some%20link.html does

not really exist, but it should show up in the RTF document without a hyperlink. It should still have a link in the PDF file created using pdflatex.

3 About paragraph breaks

There should be a paragraph break between these two urls.

   http://www.tug.org/one.
   http://www.tug.org/two.